Use the appropriate annuity formulas in this module to answer question

You are a financial adviser. Your client is thinking of investing $600 at the end of every six months for the next six years with the invested funds earning 6.4% compounded semi-annually. Your client wants to know how much money she will have after six years. What do you tell your client?
Answer #1

\large Future Value Annuity Due = Periodic Payment *\frac{(1+r)^{n}-1}{r}

where r is the rate of Return for compounding period = 6.4% /2 = 3.2% _ _ _ (Semi Annual)

n is the no of compounding period i.e 6 years * 2 = 12 periods

\large Future Value Annuity Due = 600 *\frac{(1+0.032)^{12}-1}{0.032}

=600* 14.35

= $ 8610  

After 6 years she will have $ 8610.

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.